Refer to the table. Which of the following amounts of time doing laundry would maximize net benefit?

Hours spent doing laundry 0 1 2 3 4 5
Total benefit 0 15 20 25 (a) 28
Marginal benefit 0 15 5 5 5 (b)
Total cost 0 2 5 (c) 14 (d)
Marginal cost 0 2 3 4 (e) 6

Group of answer choices

1 hour

5 hours

2 hours

3 hours
Answer #1

Option (D).

When hours = 4, Total benefit = 25 + 5 = 30 (value of a)

When hours = 3, Total cost = 5 + 4 = 9 (value of c)

When hours = 4, Marginal cost = Change in TC / Change in hours = (14 - 9) / (4 - 3) = 5/1 = 5 (value of e)

When hours = 4, Total cost = 14 + 5 = 19 (value of d)

Net benefit (NB) = Total benefit (TB) - Total cost (TC)

Hours TB TC NB
0 0 0 0
1 15 2 13
2 20 5 15
3 25 9 16
4 30 14 16
5 28 19 9

Net benefit is maximized when hours = 3 and 4.